# TapNav: Adaptive Spatiotactile Screen Readers for Tactually Guided Touchscreen Interactions for Blind and Low Vision People

### v1

@@ 1. Introduction
For example, imagine a BLV renter is exploring a scatter plot of apartment size versus rent in New York City. The screen reader speaks “Manhattan, 33rd St, 500 ft², \$3,200”; “Brooklyn, 55 Clark St, 650 ft², \$2,100”; and so on for multiple points. Although each data point is accessible, linear narration makes it challenging perceiving patterns like clusters of affordable, larger apartments. Similar challenges emerge in everyday applications with table structures (Jiang et al., 2023), such as navigating banking transactions. In these applications users must traverse entries sequentially, making it difficult to identify spending patterns. In short, audio alone conveys values but not the spatial relationships for fully understanding data representations (Munzner, 2025; Jung et al., 2021), exploration, and interface navigation that rely on position as a means to encode information and define interaction (Munzner, 2025).
@@
However, designing spatiotactile screen readers that can adapt across use cases presents several challenges. First, a tactile overlay must accommodate different layouts while maintaining consistent spatial mappings. Second, the system feedback must be designed in such way that it meaningfully binds the tactile cues to digital content without overwhelming the user. Third, designers must balance simplicity and specificity needed for effective task support (e.g., accommodate general-purpose versus task specific).
@@
  • We provide insights from Blind and Low Vision people highlighting how touch points and spatially constrained navigation can improve exploration efficiency by shifting some auditory information to tactile perception. This approach helps BLV users anticipate interaction outcomes (e.g., “The first marker should give me information about the movies with lowest ratings”) to reduce cognitive load and increasing navigation speed.
@@ 2.1. Audio-based Approaches for Touchscreen Accessibility
However, this linear narration “flattens” the interface, losing both spatial layout and hierarchical spatial relationships of screen elements that sighted people can perceive at a glance. For BLV people, accessing spatially separated points on a screen is difficult because touchscreen devices smooth surface lacks any differentiable tactile qualities (Kane et al., 2008): a point close to a corner feels the same way as a point in the middle of the screen. Thus, while navigating an interface, BLV people are forced to hold lists of screen elements in working memory and to reconstruct spatial relationships mentally (Wijntjes et al., 2008). This cognitive burden is exacerbated when using applications with interfaces designed primarily for sighted people (McGookin et al., 2008; Kane et al., 2008). To mitigate these shortcomings, researchers have explored enhancing the auditory channel in two main ways.
@@ 2.1.3. Limitations of Audio-based Approaches
While audio menus, spatialized cues, and sonification have made touchscreen content more accessible, these approaches have limitations. Audio interfaces can only convey small amounts of spatial information at once, requiring sequences of actions for both exploration and input (Palani and Giudice, 2017). This creates a load on users’ working memory (Brown et al., 2003), and even with accessibility aids, BLV users navigation with screen readers remains slow and inaccurate with complex graphical information (Sharif et al., 2021)
@@ 2.2. Touch-based Approaches for Touchscreen Accessibility
Touch can conveys information such as shapes, boundaries, and spatial configurations in parallel with audio, easing the working-memory demands that arise with audio-only solutions. To leverage this strength on touchscreens, researchers have explored three main strategies: (1) physicalizing UI controls through hardware add-ons, (2) creating tactile overlays that augment screen readers, and (3) using vibrotactile feedback to mark positions and screen elements. We refer to approaches 2 and 3 as spatiotactile screen readers—systems that provide physical feedback through tactile overlays or vibrotactile output, where each physical element corresponds to specific digital content. We review these three strategies in the following subsections.
@@ 2.2.2. Single-Layout Tactile Overlays
Raised-line, cutouts and textured sheets have long enabled BLV users to explore maps, diagrams, and other spatial content through touch (Wall and Brewster, 2006). To improve touchscreen accessbility, this idea was implemented by overlaying fixed plates that align with on-screen elements (Kane et al., 2013; Soviak et al., 2015). For example, Touchplates (Kane et al., 2013) let users overlay plastic cutouts of physical keyboards, numpads, or map guides onto a large display and reported faster target-acquisition times and higher confidence than audio-only exploration. Other researchers designed one-to-one overlays to help users understand other kinds of layouts such as document structure (Avila et al., 2018), STEM education diagrams (Melfi et al., 2020; Maćkowski et al., 2023a) and floor-plans (He et al., 2017).
@@ 3. Formative Studies
To ensure our work was adequately grounded in real needs and use cases for BLV people, we conducted two formative studies. First, a pilot test to inform the feasibility and usability of combining tactile overlays with spatial screen reader interactions for touch screen content (adaptive spatiotactile screen readers). Second, a focus groups with professionals who have experience working with BLV people to complement the technical insights from the pilot with real-world application scenarios for our main study.
@@ 3.1. Pilot Study
As a result, we enhanced our proof of concept with several new features and explored additional designs for the main user study. For example, we developed the spatial navigation mode in the interface navigation prototype (See Section 4.3) to support quick location-based access to screen elements positioned in the middle of the screen. We also introduced Braille raised markers as an alternative design to the cutout markers to leverage BLV users prior experience with this medium (See Section 4.1.1).
@@ 3.2. Focus Groups
  • Customer-facing agents face challenges when mediating access to visual information and providing guidance to BLV people using accessibility tools. While agents can describe what appears on screen or answer questions about digital documents, the mismatch between their experience using touchscreen and BLV people makes it difficult for agents to give navigation instructions that align with how BLV people actually experience digital interfaces when using their own devices (e.g., an agent might say “click the red button in the top right corner” when the screen reader announces it as “submit button, link” in a different reading order).
@@ 4.1.2. Geometric Shape Cutout Markers Overlay
Figure 4. Both exploration mode (on the left) and spatial navigation mode (on the right) in the interface navigation prototype used similar gestures as the Data Visualization prototype but included new swiping gestures to control navigation. The spatial navigation mode also showcases how participant select specific interaction areas by touching a physical marker to constrain their navigation inputs. In this example, when the user touches the Coffee Shop label, the screen reader does not provide any auditory feedback because it is outside the currently selected screen division of the Braille letter K (Column 11). If the user wants to access this label, they would either select Row ”D” or Column ”F” which have line of sight of this label, or they can go on exploration mode.:
@@ 6.1. Tactile Overlays
In this section, we present participants insights to understand how to help them better localize content and perform movements with the guidance of tactile cues, and additional considerations to inform the design of tactile markers (e.g., Low Vision versus Blind users needs, perception challenges, etc.).
@@ 6.1.3. Tactile Perception Challenges and Design Recommendations for Tactile Markers
For the Braille markers, participants discussed the Braille cells’ orientation and their expectations for a grid-based system. P3 explained that reading Braille vertically is difficult because they “appear different” when dragging their finger from top to bottom: “Like, you have to kind of switch your thinking (…) is it an M or an U? It’s an U but it’s sideways.” While only P3 explicitly mentioned the added cognitive load of reading Braille vertically (e.g., turned 90 degrees clockwise), other participants like P8 turned around the device sometimes to read the Braille markers horizontally.
@@ 6.3.3. Content, and Navigable Structures
While row and column summaries provided a clear overview of content, participants highlighted the need for complementary navigation tools like search functions and topic-based grouping to enhance both exploration and targeted access. These findings underscore the importance of layered navigation strategies, which we further explore section 7.

### v2

@@ 1. Introduction
For example, imagine a BLV renter is exploring a scatter plot of apartment size versus rent in New York City. The screen reader speaks “Manhattan, 33rd St, 500 ft², \$3,200”; “Brooklyn, 55 Clark St, 650 ft², \$2,100”; and so on for multiple points. Although each data point is accessible, linear narration makes it challenging to perceive patterns like clusters of affordable, larger apartments. Similar challenges emerge in everyday applications with table structures (Jiang et al., 2023), such as navigating banking transactions. In these applications users must traverse entries sequentially, making it difficult to identify spending patterns. In short, audio alone conveys values but not the spatial relationships for fully understanding data representations (Munzner, 2025; Jung et al., 2021), exploration, and interface navigation that rely on position as a means to encode information and define interaction (Munzner, 2025).
@@
However, designing spatiotactile screen readers that can adapt across use cases presents several challenges. First, a tactile overlay must accommodate different layouts while maintaining consistent spatial mappings. Second, the system feedback must be designed in such a way that it meaningfully binds the tactile cues to digital content without overwhelming the user. Third, designers must balance simplicity and specificity needed for effective task support (e.g., accommodate general-purpose versus task specific).
@@
  • We provide insights from Blind and Low Vision people highlighting how touch points and spatially constrained navigation can improve exploration efficiency by shifting some auditory information to tactile perception. This approach helps BLV users anticipate interaction outcomes (e.g., “The first marker should give me information about the movies with lowest ratings”) to reduce cognitive load and increase navigation speed.
@@ 2.1. Audio-based Approaches for Touchscreen Accessibility
However, this linear narration “flattens” the interface, losing both spatial layout and hierarchical spatial relationships of screen elements that sighted people can perceive at a glance. For BLV people, accessing spatially separated points on a screen is difficult because touchscreen devices’ smooth surface lacks any differentiable tactile qualities (Kane et al., 2008): a point close to a corner feels the same way as a point in the middle of the screen. Thus, while navigating an interface, BLV people are forced to hold lists of screen elements in working memory and to reconstruct spatial relationships mentally (Wijntjes et al., 2008). This cognitive burden is exacerbated when using applications with interfaces designed primarily for sighted people (McGookin et al., 2008; Kane et al., 2008). To mitigate these shortcomings, researchers have explored enhancing the auditory channel in two main ways.
@@ 2.1.3. Limitations of Audio-based Approaches
While audio menus, spatialized cues, and sonification have made touchscreen content more accessible, these approaches have limitations. Audio interfaces can only convey small amounts of spatial information at once, requiring sequences of actions for both exploration and input (Palani and Giudice, 2017). This creates a load on users’ working memory (Brown et al., 2003), and even with accessibility aids, BLV users’ navigation with screen readers remains slow and inaccurate with complex graphical information (Sharif et al., 2021)
@@ 2.2. Touch-based Approaches for Touchscreen Accessibility
Touch can convey information such as shapes, boundaries, and spatial configurations in parallel with audio, easing the working-memory demands that arise with audio-only solutions. To leverage this strength on touchscreens, researchers have explored three main strategies: (1) physicalizing UI controls through hardware add-ons, (2) creating tactile overlays that augment screen readers, and (3) using vibrotactile feedback to mark positions and screen elements. We refer to approaches 2 and 3 as spatiotactile screen readers—systems that provide physical feedback through tactile overlays or vibrotactile output, where each physical element corresponds to specific digital content. We review these three strategies in the following subsections.
@@ 2.2.2. Single-Layout Tactile Overlays
Raised-line, cutouts and textured sheets have long enabled BLV users to explore maps, diagrams, and other spatial content through touch (Wall and Brewster, 2006). To improve touchscreen accessibility, this idea was implemented by overlaying fixed plates that align with on-screen elements (Kane et al., 2013; Soviak et al., 2015). For example, Touchplates (Kane et al., 2013) let users overlay plastic cutouts of physical keyboards, numpads, or map guides onto a large display and reported faster target-acquisition times and higher confidence than audio-only exploration. Other researchers designed one-to-one overlays to help users understand other kinds of layouts such as document structure (Avila et al., 2018), STEM education diagrams (Melfi et al., 2020; Maćkowski et al., 2023a) and floor-plans (He et al., 2017).
@@ 3. Formative Studies
To ensure our work was adequately grounded in real needs and use cases for BLV people, we conducted two formative studies. First, a pilot test to inform the feasibility and usability of combining tactile overlays with spatial screen reader interactions for touch screen content (adaptive spatiotactile screen readers). Second, a focus group study with professionals who have experience working with BLV people to complement the technical insights from the pilot with real-world application scenarios for our main study.
@@ 3.1. Pilot Study
As a result, we enhanced our proof of concept with several new features and explored additional designs for the main user study. For example, we developed the spatial navigation mode in the interface navigation prototype (See Section 4.3) to support quick location-based access to screen elements positioned in the middle of the screen. We also introduced Braille raised markers as an alternative design to the cutout markers to leverage BLV users’ prior experience with this medium (See Section 4.1.1).
@@ 3.2. Focus Groups
  • Customer-facing agents face challenges when mediating access to visual information and providing guidance to BLV people using accessibility tools. While agents can describe what appears on screen or answer questions about digital documents, the mismatch between their experience using touchscreens and BLV people’s makes it difficult for agents to give navigation instructions that align with how BLV people actually experience digital interfaces when using their own devices (e.g., an agent might say “click the red button in the top right corner” when the screen reader announces it as “submit button, link” in a different reading order).
@@ 4.1.2. Geometric Shape Cutout Markers Overlay
Figure 4. Both exploration mode (on the left) and spatial navigation mode (on the right) in the interface navigation prototype used similar gestures as the Data Visualization prototype but included new swiping gestures to control navigation. The spatial navigation mode also showcases how participants select specific interaction areas by touching a physical marker to constrain their navigation inputs. In this example, when the user touches the Coffee Shop label, the screen reader does not provide any auditory feedback because it is outside the currently selected screen division of the Braille letter K (Column 11). If the user wants to access this label, they would either select Row ”D” or Column ”F” which have line of sight of this label, or they can go on exploration mode.:
@@ 6.1. Tactile Overlays
In this section, we present participants’ insights to understand how to help them better localize content and perform movements with the guidance of tactile cues, and additional considerations to inform the design of tactile markers (e.g., Low Vision versus Blind users needs, perception challenges, etc.).
@@ 6.1.3. Tactile Perception Challenges and Design Recommendations for Tactile Markers
For the Braille markers, participants discussed the Braille cells’ orientation and their expectations for a grid-based system. P3 explained that reading Braille vertically is difficult because they “appear different” when dragging their finger from top to bottom: “Like, you have to kind of switch your thinking (…) is it an M or an U? It’s a U but it’s sideways.” While only P3 explicitly mentioned the added cognitive load of reading Braille vertically (e.g., turned 90 degrees clockwise), other participants like P8 turned around the device sometimes to read the Braille markers horizontally.
@@ 6.3.3. Content, and Navigable Structures
While row and column summaries provided a clear overview of content, participants highlighted the need for complementary navigation tools like search functions and topic-based grouping to enhance both exploration and targeted access. These findings underscore the importance of layered navigation strategies, which we further explore in section 7.